# Postmerger: a new and dominant contribution to the gravitational-wave background from binary neutron stars

### v1

@@ III Results
Black solid and dashed lines in Fig. 3 represent the power-law integrated sensitivity curves of two 3G detector configurations (see Appendix C), namely, a single ET configuration (dashed lines) and an admittedly optimistic configuration consisting of two ET detectors, as well as two CE detectors (solid line), such type of network can drastically improve the background sensitivity [81, 82, 83]. Sensitivity curves for both scenarios are shown either for one-year integration time (grey solid lines) or for ten years of observations (black lines) and with an SNR for SGWB detection of $\rho=1$ [see Eq. (13)].
@@ IV Conclusion
The SGWB generated by the inspiral and merger of binary neutron stars has so far been modelled assuming that the inspiral is promptly followed by the collapse of the merger remnant to a rotating black hole. While this is reasonable for very massive binaries, it is not what is expected to take place for systems having masses around the peak of the mass distribution of neutron-star binaries. Hence, the most likely outcome of the merger is that of producing a remnant that will survive for hundreds of milliseconds before collapse (in the case of the GW170817 event, the merger lifetime has been estimated to be of about one second [84, 85]). During its short existence, the remnant can radiate enormous amounts of energy in GWs and, indeed, the radiated GW energy is 3-4 times larger than that emitted over the long inspiral.
@@
These considerations have motivated us to include also the postmerger contribution to the PSD, and hence calculate how the SGWB is modified by this important and dominant contribution. To this scope, we have implemented an analytic model of the postmerger GW signal that has been shown to provide a reasonably accurate description of the signal computed by numerical-relativity simulations [62]. Over the timescales considered here, this analytic waveform is dominated by the fundamental frequency of oscillation of the remnant $f_{2}/2$, that provides a narrow-banded emission around $f_{2}\simeq 1-2\,{\rm kHz}$. Quasi-universal relations have then been employed to relate the fundamental frequency with the tidal deformability extracted from a large sample of $10^{3}$ EOSs built in a parametric way and satisfying all the known astrophysical and microphysical constraints. Additional quasi-universal relations have also been employed to relate the tidal deformability to the chirp mass of the binary. Finally, the analytic description of the GW emission and the parametric construction of the EOSs have then been combined with a model for the population of equal-mass BNS binaries [55]. In this way, the postmerger contribution depends mostly on only two parameters: the decay timescale of the GW at the $f_{2}$ frequency, i.e., $\tau_{2}$, and the threshold value of the SNR at which a BNS can be resolved out of the background, i.e., ${\rm SNR_{thr}}$. Because both $\tau_{2}$ and ${\rm SNR_{thr}}$ are poorly known, in our analysis they are varied across the reasonable intervals in which they are expected to range.
@@
Using this setup, we were able to compute the high-frequency postmerger contribution to the SGWB from about $\sim 3.8\times 10^{5}$ binaries (i.e., those that are expected to merge in the whole universe in one year [55]), finding that this indeed represents a new and dominant contribution. While being broad, this contribution peaks at $\sim 1-2\,{\rm kHz}$ and provides a contribution to the normalised GW energy density $\Omega_{\rm{GW}}\simeq 10^{-10}-10^{-9}$, hence much larger than what is expected when considering scenarios with the inspiral and collapse only. Clearly, being a background signal, the detectability of this contribution will also depend on the integration time employed and our analysis has shown that long integration times and a network of 3G detectors are likely to be needed for the detection. However, such a network setup could detect the peak of the high-frequency component in less than one year if $\tau_{2}\gtrsim 100\,{\rm ms}$. More importantly, the detection of the high-frequency portion of the SGWB can provide precious information on the lifetime of the remnant and hence on the EOS. Overall, the results presented in this first analysis show that the postmerger contribution to the SGWB cannot be ignored as done so far, and that it contains crucial information on the HMNS. It is therefore important to extend the modelling of the SGWB to higher frequencies and to consider the detection of this background with 3G observatories.
@@
Notwithstanding the innovative nature of the procedure presented here, our approach can be improved in a number of ways. First, by considering that the decay time is not a simple mass dependant power-law, but actually a function of the EOS. Second, by extending the analysis to binaries that are not symmetric and hence introduce a degeneracy in the modelling of the chirp mass. Finally, by considering a more careful analysis of the foreground subtraction, which has been here treated simply in terms of the threshold SNR. While we do not expect any of these improvements to change the qualitative aspects of our results, they could impact the quantitative ones. We will explore these refinements of the model in future work.
V Acknowledgements
@@
This research is supported by the École Normale Supérieure Paris-Saclay CDSN PhD grant. This research is also supported in part by the ERC Advanced Grant “JETSET: Launching, propagation and emission of relativistic jets from binary mergers and across mass scales” (grant No. 884631), by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 “Strong-interaction matter under extreme conditions” – project number 315477589 – TRR 211, and by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006). LR acknowledges the Walter Greiner Gesellschaft zur Förderung der physikalischen Grundlagenforschung e.V. through the Carl W. Fueck Laureatus Chair.
Appendix A Modelling the BNS waveforms
A.1 Modelling the Inspiral
@@
To model the inspiral part of the GW waveform, we use the IMRPhenomD_NRTidalv2 model [17, 18] from the LALSuite simulation package [86]. This model extends the IMRPhenomD waveform [87, 88], designed for binary black hole systems, by incorporating EOS-dependent tidal effects. Specifically, the template uses the dimensionless tidal deformability parameter $\Lambda:=G^{4}\lambda/c^{5}M^{5}$, where $M$ is the stellar mass and the tidal deformability parameter $\lambda$ is defined as the ratio between the induced quadrupole and the external gravitational field ($G$ and $c$ are the gravitational constant and the speed of light, respectively). The model assumes that the merger remnant promptly collapses to a black hole, thus leading to very little power in the postmerger phase [18]. As commented above, this is a very drastic assumption and applicable only to a rather restricted set of binaries with total gravitational mass that is roughly given by $M_{\rm tot}\gtrsim 1.45\,M_{{}_{\rm TOV}}\,\sim 3.38\,M_{\odot}$ [89], where $M_{{}_{\rm TOV}}$ is the maximum nonrotating gravitational mass for a given EOS and the estimate comes from assuming a maximum mass $M_{{}_{\rm TOV}}\sim 2.33\,M_{\odot}$ [67, 68] (see also [90, 91, 92, 93] for a discussion on the threshold mass in BNS mergers).
A.2 Modelling the Postmerger
Numerical-relativity simulations suggest that BNS mergers typically produce a metastable HMNS, which persists for several tens of milliseconds before ultimately collapsing into a spinning black hole [22, 94]. Recent studies showed that the GW signal emitted by this HMNS exhibits several prominent peaks in the frequency domain [24, 25, 27, 28, 26, 29, 30].
Here, we use the results of Ref. [62], which presents analytical fits for the postmerger GW signal based on an extensive catalogue of time-domain numerical waveforms. These fits express the time-domain strain as a sum of damped sinusoids with a time-evolving instantaneous frequency as follows:
$h_{+}(t)=h_{0}\,\mathrm{e}^{-t/\tau_{1}}\left[\mathrm{sin}(2\pi f_{1}t)+\mathrm{sin}(2\pi(f_{1}-f_{1\epsilon})t)\right.\,\\ \hskip 59.46614pt\left.+\mathrm{sin}(2\pi(f_{1}+f_{1\epsilon})t)\right]\\ +\mathrm{e}^{-t/\tau_{2}}\,\mathrm{sin}\left(2\pi f_{2}t+2\pi\gamma_{2}t^{2}+2\pi\xi_{2}t^{3}+\pi\beta_{2}\right)\,,$ (8)
@@ A.2 Modelling the Postmerger
where $h_{0}=G\mathcal{M}_{\mathrm{chirp}}/(2Dc^{2})$ with $D$ the distance to the source, $t=0$ is the merger time, the frequencies $f_{1}$ and $f_{2}$ represent the dominant postmerger GW modes with $f_{2}>f_{1}$, $\tau_{1}$ and $\tau_{2}$ are the decay times of each mode, and $f_{1\epsilon}$, $\beta_{2}$, $\gamma_{2}$ and $\xi_{2}$ are constants [24, 95, 28]. The first peak, $f_{1}$, is short-lived (with a decay time $\tau_{1}$ of the order of a few ms) and most probably originates from damped collisions of the two stellar cores immediately after the merger [96]. This peak is accompanied by a symmetric peak $f_{3}\sim 2f_{2}-f_{1}$, which we do not take into account here, as its frequency is very high and unlikely to be detected even by third-generation (3G) detectors. The second peak, $f_{2}$, corresponds to the spin frequency of the deformed HMNS and has a decay time $\tau_{2}$ of the order of several tens of milliseconds [95, 97]. However, the expected values, and mass dependencies of both $\tau_{1}$ and $\tau_{2}$ are still rather uncertain as they sensitively depend on the EOS and the microphysical effects taking place after the merger. Interestingly, when considering the information collected with the GW170817 event, some studies have constrained the maximum lifetime of the remnant from that event to approximately one second [84, 85]. Overall, it is clear that the lifetime of the HMNS can vary significantly between a few to several thousands of milliseconds.
Here, we parametrize both decay times as power laws of the form
@@
where $\mathcal{M}_{\mathrm{chirp}}:=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$ is the chirp mass and $m_{1}$, $m_{2}$ are the component masses. Note that $\gamma<0$, as higher chirp masses correspond to a more rapid collapse, and reference decay times are taken to be $\tau_{1}=5\,{\rm ms}$ and $\tau_{2}=50\,{\rm ms}$ which correspond to typical lifetimes values Ref. [62]. Because a representative value of $\tau_{2}$ is only poorly known and the value adopted above is quite conservative, we will consider later on how our results are affected when considering larger values of $\tau_{2}$. Finally, as a reference in our analysis we take the value inferred for GW170817, i.e., $\mathcal{M}_{\mathrm{chirp,ref}}=1.186\,M_{\odot}$, and set $\gamma=-3$; alternative parametrizations are discussed in Section III, where we evaluate the impact of varying $\tau_{2}$.
The parameter $f_{1\epsilon}=50\,{\rm Hz}$ introduces a small frequency modulation in the $f_{1}$ mode, while $\beta_{2}$ adjusts the temporal phase of the waveform to ensure smooth continuity between inspiral and postmerger phases. In this work we set $\beta_{2}=0$ for simplicity and adjust the phase in the frequency domain as explained below. Furthermore, higher-order contributions to the waveform modelled by Eq. (8) come from the coefficients $\gamma_{2}$ and $\xi_{2}$, but these have a relatively minor impact on the total GW power. We take these parameters to be constant and equal to their mean values over the four EOSs studied by Ref. [62], thus $\gamma_{2}=-554\,{\rm Hz}^{2}$ and $\xi_{2}=3.36\times 10^{4}\,{\rm Hz}^{3}$.
Appendix B Modelling the BNS population
@@
In order to calculate the SGWB from the entire population of unresolved BNS mergers, we need to model the merger rate as a function of redshift and system parameters. Here, we consider (equal-mass) symmetric non-eccentric binaries, so that each BNS is characterized by the component masses $m_{1}=m_{2}=m$, tidal deformability parameters $\Lambda_{1}=\Lambda_{2}=\Lambda$, redshift $z$, sky position (right ascension $\alpha$ and declination $\delta$), orbital inclination angle $\iota$, polarization angle ($\psi$), coalescence phase $\Phi_{\mathrm{c}}$ and coalescence time set to be $t_{\mathrm{c}}=0$. The mass distribution we used is derived from Galactic observations and is assumed to be independent of redshift. According to Ref. [98], this distribution is well-modeled by a Gaussian with a mean of $1.33\,M_{\odot}$ and a standard deviation of $0.09\,M_{\odot}$. Both neutron stars in the binary are assumed to follow this distribution, which is tightly concentrated around the mean, making it a non-critical parameter for this study.
@@
We note that this mass distribution may depend on the formation scenario of the binary. More specifically, the mass distribution of BNS detected via GW may differ from that of Galactic binaries, potentially explaining the higher total mass observed in the BNS merger GW190425 [11, 99]. Furthermore, the recent GW190814 event suggests a black hole-neutron star (BHNS) merger involving a neutron star significantly heavier than typical Galactic counterparts [100], although there are arguments favouring a black-hole origin for the secondary [71].
We assume that the BNS merger rate follows the star formation rate (SFR) convolved with a distribution of delay times. Specifically, we adopt the baseline_delays model from Ref. [55], where the merger rate is defined as:
$R_{\mathrm{merg}}(t)=R_{0}\int^{t_{\mathrm{d},\mathrm{max}}}_{t_{\mathrm{d},\mathrm{min}}}\mathrm{SFR}(t-t_{\mathrm{d}})\,P(t_{\mathrm{d}})\,\mathrm{d}t_{\mathrm{d}}\,.$ (10)

### v2

@@ Abstract
Black solid and dashed lines in Fig. 1 represent the power-law integrated sensitivity curves of two 3G detector configurations: a single ET configuration (dashed lines) and an admittedly optimistic configuration consisting of $2$ ET detectors, as well as two CE detectors (solid line), such type of network can drastically improve the background sensitivity Gupta, Ish et al. (2024); Reitze, David et al. (2019a); Maggiore et al. (2024). Sensitivity curves for both scenarios are shown either for one-year integration time (grey solid lines) or for ten years of observations (black lines) and with an SNR for SGWB detection of $\rho=1$.
@@
Conclusion— The SGWB generated by the inspiral and merger of binary neutron stars has so far been modelled assuming that the inspiral is promptly followed by the collapse of the merger remnant to a rotating black hole. While this is reasonable for very massive binaries, it is not what is expected to take place for systems having masses around the peak of the mass distribution of neutron-star binaries. Hence, the most likely outcome of the merger is that of producing a remnant that will survive for hundreds of milliseconds before collapse. During its short existence, the remnant can radiate enormous amounts of energy in GWs and, indeed, the radiated GW energy is 3-4 times larger than that emitted over the long inspiral.
@@
These considerations have motivated us to include also the postmerger contribution to the PSD and hence calculate how the SGWB is modified by this important and dominant contribution. To this scope, we have implemented an analytic model of the postmerger GW signal that has been shown to provide a reasonably accurate description of the signal computed by numerical-relativity simulations Bose et al. (2018). The analytic waveform is dominated by the fundamental frequency of oscillation of the remnant $f_{2}/2$, that provides a narrow-banded emission around $f_{2}\simeq 1-2\,{\rm kHz}$. Quasi-universal relations have been employed to relate $f_{2}$ with the tidal deformability and the latter with the chirp mass of the binary. Finally, the analytic description of the GW emission and the parametric construction of the EOSs have then been combined with a model for the population of equal-mass BNS binaries Lehoucq et al. (2023). In this way, the postmerger contribution depends mostly on only two parameters: the decay timescale of the GW at the $f_{2}$ frequency, i.e., $\tau_{2}$, and the threshold value of the SNR at which a BNS can be resolved out of the background, i.e., ${\rm SNR_{thr}}$. Because both $\tau_{2}$ and ${\rm SNR_{thr}}$ are poorly known, in our analysis they are varied across the reasonable intervals in which they are expected to range.
@@
Using this setup, we were able to compute the high-frequency postmerger contribution to the SGWB from about $\sim 3.8\times 10^{5}$ binaries (i.e., those that are expected to merge in the whole universe in one year Lehoucq et al. (2023)), finding that this indeed represents a new and dominant contribution. While being broad, this contribution peaks at $\sim 1-2\,{\rm kHz}$ and provides a contribution to the normalised GW energy density $\Omega_{\rm{GW}}\simeq 10^{-10}-10^{-9}$, hence much larger than what is expected when considering scenarios with the inspiral and collapse only. Clearly, being a background signal, the detectability of this contribution will also depend on the integration time employed and our analysis has shown that long integration times and a network of 3G detectors are likely to be needed for the detection. However, such a network setup could detect the peak of the high-frequency component in less than one year if $\tau_{2}\gtrsim 100\,{\rm ms}$. More importantly, the detection of the high-frequency portion of the SGWB can provide information that is either complementary or supplementary to that of the foreground signals. More specifically, the detection of the high-frequency component would constrain the typical lifetime of the remnant and since the latter depends on the EOS, it would provide complementary information on the EOS. On the other hand, when considering the reconstruction of the BNS population and its evolution with redshift, the detection of the high-frequency component – and specifically the location of the maximum and its broadness – will tell us about the whole population of binaries beyond what can be deduced from the loudest foreground sources. Overall, the results presented in this first analysis show that the postmerger contribution to the SGWB cannot be ignored as done so far and that it contains crucial information on the HMNS accessible with 3G detectors.
@@
Notwithstanding the innovative nature of the procedure presented here, our approach can be improved in a number of ways. First, by considering that the decay time is not a simple mass dependant power-law, but actually a function of the EOS. Second, by extending the analysis to binaries that are not symmetric and hence introduce a degeneracy in the modelling of the chirp mass. Finally, by considering a more careful analysis of the foreground subtraction, which has been here treated simply in terms of the threshold SNR. We will explore these refinements of the model in future work.
@@
Acknowledgements— This research is supported by the École Normale Supérieure Paris-Saclay CDSN PhD grant. This research is also supported in part by the ERC Advanced Grant “JETSET: Launching, propagation and emission of relativistic jets from binary mergers and across mass scales” (grant No. 884631), by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 “Strong-interaction matter under extreme conditions” – project number 315477589 – TRR 211, and by the State of Hesse within the Research Cluster ELEMENTS (Project ID 500/10.006). LR acknowledges the Walter Greiner Gesellschaft zur Förderung der physikalischen Grundlagenforschung e.V. through the Carl W. Fueck Laureatus Chair.
End Matter
On the energy radiated in GWs— A major focus of our work is to extend the contribution of the BNS merger signal to the SGWB by considering not only the inspiral part but also the postmerger emission. To appreciate simply why this is important and hence the relevance of our findings, we show in Fig. 3 the total GW energy radiated by a population of BNS when integrated over the entire frequency range ($E_{\mathrm{GW}}$) and normalized by the total energy considering only the inspiral phase ($E_{\mathrm{GW},0}$), as a function of the time before and after the merger $t-t_{\rm mer}$. Near the merger ($t-t_{\rm mer}\lesssim 0$) the energy radiated increases very rapidly, gaining an order of magnitude in the last $10\,{\rm ms}$ before the merger. Following the merger ($t-t_{\rm mer}>0$), the energy radiated continues to grow at a slower rate but more intensely, so that it can reach values much larger than those lost during the whole inspiral. Stated differently, BNS mergers are copious emitters of GWs during the inspiral, but it is the postmerger that provides the largest contribution to the radiated energy.
Of course, the exact amount of energy radiated will depend sensitively on the EOS describing the matter of which the BNS is composed of and which is essentially unknown. Instead of concentrating on a few specific EOSs, it is more convenient to consider a large ensemble of agnostic EOSs satisfying astrophysical, GW, or elementary-particle constraints Annala et al. (2018); Most et al. (2018); Greif et al. (2019); Altiparmak et al. (2022). This is what is represented with the blue-shaded band in Fig. 3, which reports the 99th percentile contour of an ensemble of $10^{3}$ EOSs generated using our EOS-agnostic approach. Similarly, the solid blue line reports the median of the distribution at each time in the evolution. In this way, it is straightforward to appreciate that since most of the energy is lost after the merger, it is important to include also the postmerger emission when calculating the BNS GW background (similar considerations apply when considering the radiated angular momentum Ecker et al. (2025); Bernuzzi et al. (2016)).
Figure 3: Evolution of the radiated GW energy when normalized to the at the merger. Shown with blue- and red-shaded areas are the corresponding values during the inspiral (blue) and the postmerger. The variance in the shaded regions marks the $99\%$ percentile.
@@
Modelling the BNS waveforms— To model the inspiral part of the GW waveform, we use the IMRPhenomD_NRTidalv2 model Dietrich, Tim et al. (2019); Dietrich et al. (2019) from the LALSuite simulation package LIGO Scientific Collaboration (2020). This model extends the IMRPhenomD waveform Husa et al. (2016); Khan et al. (2016), designed for binary black hole systems, by incorporating EOS-dependent tidal effects. Specifically, the template uses the dimensionless tidal deformability parameter $\Lambda:=G^{4}\lambda/c^{5}M^{5}$, where $M$ is the stellar mass and the tidal deformability parameter $\lambda$ is defined as the ratio between the induced quadrupole and the external gravitational field ($G$ and $c$ are the gravitational constant and the speed of light, respectively). The model assumes that the merger remnant promptly collapses to a black hole, thus leading to very little power in the postmerger phase Dietrich et al. (2019). As commented above, this is a very drastic assumption and applicable only to a rather restricted set of binaries with total gravitational mass that is roughly given by $M_{\rm tot}\gtrsim 1.45\,M_{{}_{\rm TOV}}\,\sim 3.38\,M_{\odot}$ Koeppel et al. (2019), where $M_{{}_{\rm TOV}}$ is the maximum nonrotating gravitational mass for a given EOS and the estimate comes from assuming a maximum mass $M_{{}_{\rm TOV}}\sim 2.33\,M_{\odot}$ Margalit and Metzger (2017); Rezzolla et al. (2018) (see also Bauswein et al. (2013); Tootle et al. (2021); Kölsch et al. (2022); Kashyap et al. (2022) for a discussion on the threshold mass in BNS mergers).
@@
Numerical-relativity simulations suggest that BNS mergers typically produce a metastable HMNS, which persists for several tens of milliseconds before ultimately collapsing into a spinning black hole Baiotti and Rezzolla (2017); Paschalidis (2017). Recent studies showed that the GW signal emitted by this HMNS exhibits several prominent peaks in the frequency domain Stergioulas et al. (2011); Bauswein and Janka (2012); Bauswein and Stergioulas (2015a); Rezzolla and Takami (2016); Takami et al. (2014a, 2015a); Topolski et al. (2024). Here, we use the results of Ref. Bose et al. (2018), which presents analytical fits for the postmerger GW signal based on an extensive catalogue of time-domain numerical waveforms. These fits express the time-domain strain as a sum of damped sinusoids with a time-evolving instantaneous frequency as $h_{+}(t)=h_{0}\,\mathrm{e}^{-t/\tau_{1}}[\mathrm{sin}(2\pi f_{1}t)+\mathrm{sin}(2\pi(f_{1}-f_{1\epsilon})t)+\mathrm{sin}(2\pi(f_{1}+f_{1\epsilon})t)]+\mathrm{e}^{-t/\tau_{2}}\,\mathrm{sin}(2\pi f_{2}t+2\pi\gamma_{2}t^{2}+2\pi\xi_{2}t^{3}+\pi\beta_{2})$, where $h_{0}=G\mathcal{M}_{\mathrm{chirp}}/(2Dc^{2})$ with $D$ the distance to the source, $t=0$ is the merger time, the frequencies $f_{1}$ and $f_{2}$ represent the dominant postmerger GW modes with $f_{2}>f_{1}$, $\tau_{1}$ and $\tau_{2}$ are the decay times of each mode, and $f_{1\epsilon}$, $\beta_{2}$, $\gamma_{2}$ and $\xi_{2}$ are constants Stergioulas et al. (2011); Takami et al. (2014b); Rezzolla and Takami (2016). The first peak, $f_{1}$, is short-lived (with a decay time $\tau_{1}$ of the order of a few ms) and most probably originates from damped collisions of the two stellar cores immediately after the merger Takami et al. (2015b). The second peak, $f_{2}$, corresponds to the spin frequency of the deformed HMNS and has a decay time $\tau_{2}$ of the order of several tens of milliseconds Takami et al. (2014b); Bauswein and Stergioulas (2015b). However, the expected values, and mass dependencies of both $\tau_{1}$ and $\tau_{2}$ are still rather uncertain as they sensitively depend on the EOS and the microphysical effects taking place after the merger. Interestingly, when considering the information collected with the GW170817 event, some studies have constrained the maximum lifetime of the remnant from that event to approximately one second Gill et al. (2019); Murguia-Berthier et al. (2021). Overall, it is clear that the lifetime of the HMNS can vary significantly between a few to several thousands of milliseconds.
@@
Figure 4 illustrates the construction of the hybrid waveform for a BNS system with masses $M_{1}=M_{2}=1.25\,M_{\odot}$, at a distance from observer $D=50\,{\rm Mpc}$ when employing GNH3 EOS discussed in Ref. Bose et al. (2018). Shown with a blue dashed line is the inspiral-and-merger signal followed by a prompt collapse and described by the IMRPhenomD_NRTidalv2 model, the postmerger signal with a black dashed line, and the combined (hybrid) signal with a red solid line. We also show with a solid green line the ET strain sensitivity. The transition frequency $f_{\rm tr}$ is meant to mark therefore the boundary between the grey and white regions. In the grey region, the hybrid waveform is simply given by the IMRPhenomD_NRTidalv2 waveform. Note that in the white region, the hybrid waveform is given by the sum of the inspiral and postmerger strains and that the latter is negligible below $f_{\rm tr}$, thus ensuring no significant power is lost or added. To properly combine the two waveforms above $f_{\rm tr}$, we must ensure phase coherence, which is not guaranteed a priori, as we have not explicitly computed the temporal phase $\beta_{2}$. Therefore, at $f_{\rm tr}$, we adjust the phase of the IMRPhenomD_NRTidalv2 waveform to match the postmerger one, thus ensuring phase continuity.
Figure 4: Evolution in the frequency-domain of the GW strain for a BNS with masses $m_{1}=m_{2}=1.25\,M_{\odot}$ at a distance of observer $D=50\,{\rm Mpc}$ and relative to the GNH3 EOS Bose et al. (2018). The IMRPhenomD_NRTidalv2 waveform is shown with a black dashed line, the HMNS postmerger with a blue dashed line, and the combined waveform with a red solid line. The frequency $f_{\rm tr}$ marks the transition between the inspiral (grey-shaded are) and the postmerger, while the green solid curve reports the sensitivity of ET. Also reported are the main components $f_{2}$ and $f_{1}$ of the spectrum of equal-mass binaries.
@@
Modelling the BNS population— In order to calculate the SGWB from the entire population of unresolved BNS mergers, we need to model the merger rate as a function of redshift and system parameters. Here, we consider (equal-mass) symmetric non-eccentric binaries, so that each BNS is characterized by the component masses $m_{1}=m_{2}=m$, tidal deformability parameters $\Lambda_{1}=\Lambda_{2}=\Lambda$, redshift $z$, sky position (right ascension $\alpha$ and declination $\delta$), orbital inclination angle $\iota$, polarization angle ($\psi$), coalescence phase $\Phi_{\mathrm{c}}$ and coalescence time set to be $t_{\mathrm{c}}=0$. The mass distribution we used is derived from Galactic observations and is assumed to be independent of redshift. According to Ref. Farrow et al. (2019), this distribution is well-modeled by a Gaussian with a mean of $1.33\,M_{\odot}$ and a standard deviation of $0.09\,M_{\odot}$. Both neutron stars in the binary are assumed to follow this distribution, which is tightly concentrated around the mean, making it a non-critical parameter for this study.
@@
We note that this mass distribution may depend on the formation scenario of the binary. More specifically, the mass distribution of BNS detected via GW may differ from that of Galactic binaries, potentially explaining the higher total mass observed in the BNS merger GW190425 collaboration (2020); Romero-Shaw et al. (2020). Furthermore, the recent GW190814 event suggests a black hole-neutron star (BHNS) merger involving a neutron star significantly heavier than typical Galactic counterparts Abbott, R. et al. (2020), although there are arguments favouring a black-hole origin for the secondary Nathanail et al. (2021).